OpenStudy (nuccioreggie):

What is the measure of angle x? A right angle is shown divided into two parts. The measure of one part of the right angle is 30 degrees. The measure of the other part is 3x. 10 20 30 40

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

@amistre64

OpenStudy (nuccioreggie):

@freckles

OpenStudy (anonymous):

HI!

OpenStudy (anonymous):

This question is actually really simple. You are probably just thinking about it too hard. When it's a right triangle you know one angle is 90 degrees. The other two angles are always the same with a right triangle. Since one of them is 30 degrees, then the other is also 30 degrees. To find x you must do 30/3=?. So the answer is A. x=10.

OpenStudy (nuccioreggie):

No ur wrong i already got the answer its B sorry

OpenStudy (nuccioreggie):

it is 3x=90-30 3x=60 x=20